Machines A & B have been showing disparate efficiency and defects per unit rates despite similar variation in quality at purchase time. If the mean squares between the 2 machines is 250 and the mean square from each machine is 100,what is the value of FSTAT(F-Statistics)?
Answer #1

We are given :

(among variance) MSA =250

(variance within) MSW =100

FSTAT is the ratio of variance value among and within, I. E. = MSA /MSW = 250/100=2.5